OpenStudy (anonymous):

are these all real numbers? sqt-15 44 1.765 -1.5 sqt-6 0

OpenStudy (anonymous):

some are

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

true

myininaya (myininaya):

no they are not all real \[\sqrt{-15}=i \sqrt{15}, \sqrt{-6}=i \sqrt{6}\] these two are imaginary

myininaya (myininaya):

others are real

OpenStudy (anonymous):

thank you !!!!

OpenStudy (anonymous):

is sqt-16 real?